OpenStudy (anonymous):

2√6 + 3√96 (8+√11)(8-√11) 4=√p -2 √q-9 =7 the rad sign is over the q and -9 √5x-1 =√4x+9 √p=-1 # –1 is a solution of the original equation. 1 is an extraneous solution. # 1 is a solution of the original equation. # 1 is a solution of the original equation. –1 is an extraneous solution. # There is no solution. y=√3x+3 # x ≤ – 1 # x > 1 # x ≥ –1

OpenStudy (anonymous):

are there 7 questions.. or you showed your work !! that you attempted !!

OpenStudy (anonymous):

no 7 questions i dont get these ones at all

OpenStudy (anonymous):

what don't you get? have you tried at all ? ... where are you stuck, Let me know so I can help you ! :S

OpenStudy (anonymous):

i dont understand how to solve any of these or how to begin to find the answers its on my homework and ive completed all the rest besides these ones

OpenStudy (anonymous):

\[\Large 2\sqrt{6}+3\sqrt{96}=2\sqrt{6}+3\sqrt{16\cdot 6}\] \[\Large 2\sqrt6+3\cdot 4\sqrt6 =2\sqrt6+12\sqrt6\] now sum them... and you'll have \[\Large 14\sqrt6\] This was a hint regarding to the first question

OpenStudy (anonymous):

the second one .. just multiply, and tell me if you'll get stuck, so I can give a hint again... You'll forget these kinds of problems very soon , if you don't try them yourself !! ;)

OpenStudy (anonymous):

im not understanding the second one though like i see how you did the first but the second is different what do i multiply

OpenStudy (anonymous):

\[\Large (8+\sqrt{11})(8-\sqrt{11})=(8\cdot 8)-8\sqrt{11}+8\sqrt{11}-(\sqrt{11})^2\]

OpenStudy (anonymous):

64-11=53 is the answer ;) ... NOTE: \[\Large -8\sqrt{11}+8\sqrt{11}=0\] so they cancel out ! :) try the next one ;)

OpenStudy (anonymous):

ok so is the next one 36? i think

OpenStudy (anonymous):

how do you get there .. show your work ! :P LOL

OpenStudy (anonymous):

..never mind, that's correct ;) well done ! \[\Large 4=\sqrt p-2 \longrightarrow \sqrt p=4+2\] now square both sides ... \[\Large p=36\] If this is what you've done... that's correct ;) .. try the next one

OpenStudy (anonymous):

ok yay! let me try the next one

OpenStudy (anonymous):

is it 58?

OpenStudy (anonymous):

58-9= 49 squared is 7?

OpenStudy (anonymous):

\[\Large \sqrt{q-9}=7 \longrightarrow q-9=49\] q=49+9 q=58

OpenStudy (anonymous):

is that what you got ?

OpenStudy (anonymous):

I mean by calculating .. the result is it? 1=58 ?

OpenStudy (anonymous):

q=58

OpenStudy (anonymous):

yah thats what im saying

OpenStudy (anonymous):

.. then that's correct... y=√3x+3 # x ≤ – 1 # x > 1 # x ≥ –1 \[\large y=\sqrt{3x+3}\] we have...\[\large \sqrt{3x+3}\geq 0\] square both sides... \[\Large 3x+3\geq 0\] \[\Large 3x\geq -3 \longrightarrow x\geq \frac{-3}{3}\] \[\Large x\geq -1\]

OpenStudy (anonymous):

this was the last one...

OpenStudy (anonymous):

no the √p=-1

OpenStudy (anonymous):

tell me a number who when you square it, it becomes negative ??

OpenStudy (anonymous):

√p=-1 # –1 is a solution of the original equation. 1 is an extraneous solution. # 1 is a solution of the original equation. # 1 is a solution of the original equation. –1 is an extraneous solution. # There is no solution.

OpenStudy (anonymous):

\[\Large (-1)^2 \neq -1\] so.. there's no number which when we square it.. it becomes NEGATIVE ....it doesn't exist.

OpenStudy (anonymous):

so there is no solution

OpenStudy (anonymous):

exactly ! :)

OpenStudy (anonymous):

thankyouu!!!

OpenStudy (anonymous):

wait is there anyway you could just look at his last question????

OpenStudy (anonymous):

@haileystowers sorry... the lat one, you'll do it yourself , I've to go :) ... Glad to help

OpenStudy (anonymous):

about triangle... use Tangent since \[\Large \tan x=\frac{\sin x}{\cos x}\] we have... \[\Large \frac{\sin 42^{\circ}}{\cos 42^{\circ}}=\frac x5 \] use cross multiplication and the answer should be ... \[\Large x\approx 4.5\] sorry .. got to go... use calculator or something, Have a nice day . Bye :)

OpenStudy (anonymous):

thankyou!